Solve by simplex method. Max z = 2x1 -x2 +2x3, 2x1 +x2 < 10, x1 +2x2 -2x3 <20x1 +2x3 < 5, x1, x2, x3 > 0


1
Expert's answer
2021-11-11T12:15:06-0500

After introducing slack variables

Max Z=2x1-x2+2x3+0S1+0S2+0S3

subject to

2x1+x2+S1=10x1+2x2-2x3+S2=20x1+2x3+S3=5

and x1,x2,x3,S1,S2,S3≥0


after iteration 1:

Negative minimum Zj-Cj is -2 and its column index is 1. So, the entering variable is x1.

Minimum ratio is 5 and its row index is 3. So, the leaving basis variable is S3.

∴ The pivot element is 1.

Entering =x1, Departing =S3, Key Element =1


after iteration 2:

Since all Zj-Cj≥0


Hence, optimal solution is arrived with value of variables as :

x1=5,x2=0,x3=0

Max Z=10